### v2

@@ 3.2 Stylized facts II: hierarchical positions and geographical distances
Table 4: How distant are subsidiaries from parent companies? - Complex MNEs
@@
  • Note: The table reports descriptive evidence on foreign subsidiaries controlled by multinational enterprises. The simplest MNEs with only one direct subsidiary are excluded. Columns 1, 2 and 3 on the left indicate the position in the ownership hierarchy, i.e., how many control links run from the HQs to the single subsidiary. Columns 4, 5, and 6, in the centre, count how many national borders are crossed along an ownership chain to reach a final subsidiary from the headquarters. Columns 7, 8, and 9 on the right report the geographic distance (thousands km) separating final subsidiaries from parent companies. Distance is measured between the most populated cities of hosting countries.
@@
Finally, we take a more classical approach by measuring the physical distance in kilometers between headquarters and subsidiaries, as reported in columns 7 to 9 of table 4. These data, obtained from Conte et al. (2022) (CEPII database), represent the distance between the main cities in the countries where the headquarters and subsidiaries are located. As expected, 47.3% of the subsidiaries are located within 5,000 km of the parent company, a somewhat narrower range than expected. Beyond this threshold, the distribution has a long right tail, with subsidiaries located at increasing distances from the parent. [^15] The longest observed distance in our dataset is 19,950 km, between Paraguay and Taiwan.
@@ 4 A structural model for ownership chains
Motivated by previous empirical evidence, we present a simple model to explain location strategies along ownership chains in multinational enterprises (MNEs). Inspired by Head and Ries (2008), we assume a competitive contest in which headquarters compete to invest in new subsidiary assets worldwide while incurring monitoring costs. Specifically, in the case of ownership chains, each headquarters faces a choice: it can either directly monitor a final target subsidiary or delegate monitoring to an intermediate managerial unit - an intermediate affiliate (sometimes referred to as a middleman). These intermediate affiliates are subsidiaries that are already part of the corporate structure before the auction takes place. A subsidiary is added to an ownership chain when the headquarters delegates or self-delegates monitoring, i.e., when the cost of delegation is effectively zero. In the latter case, we have the special case of direct control.
@@
Our model abstracts from production processes between headquarters and subsidiaries. While our setup does not exclude the possibility of supply chain coordination, it also applies to conglomerates and diversified business portfolios, where monitoring remains crucial for governance and managerial control.161616 To assess whether the length of ownership chains is associated with increasing distance in the production space, we ran a regression in which the hierarchical distance of an affiliate from its parent served as the dependent variable, including, among other controls, the “production distance”, i.e. $upstramness_{ij}$, a measure from Alfaro et al. (2019). While the coefficient was statistically significant, its magnitude was negligible. This suggests that an implausibly long ownership chain would be required for an economically meaningful difference in production stages between parent and subsidiary. In our data, these results suggest that ownership chains are primarily shaped by organizational considerations rather than sequential production optimization. Results are available upon request.
@@ 4.1 The inspection game
In this corporate governance game, profits arise from the value generated by the headquarters (Hq) and its existing perimeter of subsidiaries ($A$), [^18] The existence of a corporate perimeter before new acquisitions implies that $A=\sum_{f}a_{f}$, meaning that the total value is generated by both Hq and all existing subsidiaries, regardless of new acquisitions., as well as by the additional subsidiary ($b$), conditional on the effort exerted by the subsidiary’s manager ($e$). Headquarters can choose to monitor $q$ or trust $1-q$ the subsidiary. Similarly, the subsidiary’s manager can work ($1-p$) or shirk ($p$). If the subsidiary works, headquarters pays a wage $w$; however, if headquarters monitors and discovers shirking, the wage and the value of the subsidiary default to zero ($w=b=0$).
@@
Since corporate control operates in a global market, all Hqs play the same inspection game when evaluating subsidiaries that are available for acquisition. Consequently, competition ensures that the winning Hq is the one offering the highest valuation, determined by equation 2. It is important to note that, compared to the classical literature on delegation and monitoring in hierarchical structures (Holmstrom and Tirole, 1989), our framework relies on a number of simplifying assumptions regarding the contractual environment. First, we assume no information asymmetry between Hqs and intermediate affiliates (Mids), i.e., each party perfectly observes the other’s effort. Second, we exclude the possibility of side contracting, thus eliminating potential collusion between Mids and Subs. Finally, we treat the effort made by Mids as exogenous. In the next section, we examine how these assumptions shape Hq’s decision process and monitoring cost design.
@@ 4.2.2 The case of direct control
The simplest strategy for Hq is to retain full corporate control and not delegate monitoring to a Mid. In this case, Hq self-delegates, which means that delegation costs are effectively zero. More formally, in our model, direct control is treated as a valid location choice without delegation, represented as a $ikj$-th triplet where the intermediate jurisdiction ($k$) is null, i.e., $k=\emptyset$.
@@ 4.3.1 Locations of the chain
where $\mathcal{M}_{ikj}=1$ if Hq in country $i$ willing to invest in country $j$ chooses her monitoring unit in country $k$. F[•] is the joint cumulative distribution function, conditional on the value of $\epsilon_{ikj}$. By integrating $F[\bullet]$ with respect to the marginal probability density function $f(\epsilon_{ikj})$, we obtain the unconditional probability - the probability that country $k$ is chosen as the monitoring location for any given realization of $\epsilon_{ikj}$.
@@
where we can define $\mathcal{C}_{ij}=-ln[\sum\limits_{\ell}e^{-(\delta_{i\ell}+\delta_{\ell j})}]$ as the expected aggregate cost of locations $j$ for Hqs in country $i$, whatever intermediate country $k$.
@@ 4.3.2 Probability of winning the auction for corporate control
The highest valuation wins the auction, i.e., the marginal probability of a Hq in country $i$ gaining control over a Sub in country $j$ is given by the probability that the highest valuation for a given Sub in $j$ comes from an Hq located in $i$, i.e., the maximum valuation in $i$ is the highest among the maxima of all competing countries. This probability is:
@@ 6 Results
Finally, the last column of Table 6 presents the coefficient of the vector of the expected cost of corporate control ($E[C_{ikj}]$), obtained from previous estimates. The coefficient obtained corresponds to $\theta$ in equation 9, and its sign and magnitude are as expected: an increase in the cost index reduces the number of subsidiaries in country $j$ controlled by parents in country $i$.
Figure 8 shows the distribution of the expected inspection cost for each origin $i$ and destination $j$, which is a useful byproduct of our structural model.
Figure 8: Distribution of the expected inspection cost
  • Note: The figure shows the distribution of the expected cost of corporate control estimated after Eq. 8.
@@ 7 Robustness and sensitivity checks
  • Note: Ratios $CT_{k}/CT_{i}$ and $CT_{j}/CT_{k}$ control for relative advantage in corporate taxation. Ratios $LC_{k}/LC_{i}$ and $LC_{j}/LC_{k}$ control for the relative advantage in labour costs. Given the presence of zeros in the case of direct control, we scale the distances in km by using the hyperbolic sine transformation ($ln(x+\sqrt{x^{2}+1})$), which allows approximating to the natural logarithm while retaining the zeros (Bellemare and Wichman, 2020). Standard errors are double-clustered by origin and destination in parentheses (*** p$<$0.01, ** p$<$0.05, * p$<$0.1).
@@ 8 Conclusions
Building on these empirical patterns, we propose a location choice model in which parent firms compete in a global auction to acquire new subsidiaries. In formulating their bids, parents weigh the option of delegating monitoring responsibilities to an intermediate affiliate in a third country. he model yields two testable equations: the first is a trilateral equation that determines whether monitoring is delegated to an intermediate affiliate and where it is located, while the second is a bilateral equation that predicts the location of newly acquired final subsidiaries.
@@
In this Appendix, we provide an understanding of how raw shareholding data are used to extract corporate trees developed by MNEs. For further details, we refer to (Rungi et al., 2017). An MNE is made of a unique parent company and its subsidiaries. If we look at their ownership data, they represent a corporate network in the form of a hierarchical tree, where nodes are legally autonomous companies. Please consider the simplified fictional corporate tree in Figure A1. On top of the hierarchy, we find the parent company, which exerts control over subsidiaries by means of direct and indirect shareholding links. In principle, an MNE can develop much more complex patterns than the ones depicted in Figure A1, as it can include cross-holdings when shareholding relationships run two-way between companies. They can also entail ownership cycles, i.e., when the firm holds a share of its equity stakes. Given a fragmentation of shareholding, companies can indirectly connect through multiple sequences of ownership links. The solution is to start with a full ownership matrix of all firms and their shareholders in the data. Thus, a control relationship is detected iteratively on a transformation of the ownership matrix. In line with international accounting standards (OECD, 2005; UNCTAD, 2009; Eurostat, 2007), one can assume that corporate control is present when there is an absolute majority of equity stakes. Yet, one cannot exclude that lower thresholds still allow the parent company the ability to influence and contribute to the decision-making of another company.
@@
In Figure A1, we easily attribute direct control of Subsidiary 1 to the Parent because the latter holds an absolute majority of voting rights in the first. Besides this most elementary case of direct control, there are two other ways for the Parent to exert control: (i) by transitivity, when successive majority links form an ownership chain, and subsidiaries align on it, as in the case of Subsidiary 3; (ii) by consolidation, when equity is fragmented, no direct majority path can be detected from the Parent, and yet other subsidiaries hold minority links in a firm, like Subsidiary 4, which can be consolidated and allow the unique Parent to finally reach out the management of Subsidiary 4.
Ownership chains. Figure A1 shows an MNE that has two ownership chains. One that starts from the Parent and reaches out to Subsidiary 3 through Subsidiary 1. A second one that starts from the Parent and reaches out to Subsidiary 4 thanks to either Subsidiary 1 or Subsidiary 3. To solve uncertainty, we will assume that the final subsidiary is on the shortest ownership chain. If ownership chains are of equal length, then we discriminate on shares. When shares are equal, we discriminate by firm size, giving priority to the middleman that is bigger.
@@
In the following paragraphs, we introduce three naïve unstructured applications of gravity models to observe correlations, which are informative for the setup of our theoretical model, and to discuss what the analyst misses when she does not consider information on ownership chains. We argue this section is important to realize empirical problems from previous literature. In each case, we test the following general Pseudo-Poisson specification at the country level:
@@
where $N_{ij}$ counts the number of companies in country $j$ controlled by companies in country $i$. We include bilateral frictions in matrix $\mathbf{X}_{ij}$. Notably, besides classical gravity controls source from the CEPII Gravity dataset (Conte et al., 2022) $wh_{ij}$, we include a measure of overlapping working hours between countries $i$ and $j$ as a proxy of the common working schedule across time zones. See further details in Section 3. Finally, $\gamma_{i}$ and $\gamma_{j}$ capture origin and destination country fixed effects. [^22] Please note that our information on MNEs’ perimeters is cross-section. Therefore, we can focus on the between-variation and cannot observe the within-variation.
At this point, we are ready with the empirical model, but how do we treat ownership chains? We will show that each of the three following cases is wrong. Or, we could say they are some reduced forms of the real model. Let us introduce the issue starting from a fictional sample, including two kinds of MNEs most common in our data set, whose fictional structures are represented in Figure B1.
@@
  • Note: This fictional dataset includes a simple and a complex MNE. Each node represents a firm, and each directed links a majority of voting rights of one firm to another
@@
  • Notes: The figure shows three fictional cases of ownership chains included in bilateral gravity models. The legends on the right report the directed links one would include in each application. We hereby discuss the issues with each of them.
More specifically, in Figure B1, we draw one simple MNE controlling only one subsidiary, which we know is the most common case, although we also know that, on aggregate, these MNES have a small economic weight (see Figures 4 and 3). A second complex MNE on the right controls seven subsidiaries organized along four ownership chains: i) A $\rightarrow$ B; ii) A $\rightarrow$ C $\rightarrow$ E $\rightarrow$ H; iii) A $\rightarrow$ D $\rightarrow$ F; iv) A $\rightarrow$ D $\rightarrow$ G.
@@
Let us first assume that linkages are not disclosed, but she knows the set of subsidiaries. In this case, she will count the parents in country $i$ and the subsidiaries in country $j$. This is the case (a) in the fictional sample redrawn in Figure B2. We guess this is the case that most scholars have in mind when testing gravity models with firm-level data on MNEs. See, for example, Mutti and Grubert (2004), which reports an empirical analysis using data on the location choices of U.S. parent companies.
A second alternative is that ownership chains are only partially disclosed because the statistician has problems collecting them. She falsely attributes the ultimate control of a subsidiary to its immediately controlling corporate shareholder. Hence, our fictional complex MNE will appear fragmented into three smaller pieces. The information included in Eq. 10 is false. This is the second case in Figure B2, where the list of links on the right does not include $C\rightarrow E$. Indeed, full disclosure of the corporate structure of MNEs is a long-standing problem for national and international statistics offices. The necessity to identify ultimate controlling entities is required by several guidelines (OECD, 2005; UNCTAD, 2009; IFRS, 2011), but seldom accomplished.
A third alternative is that the analyst knows the entire ownership chain but does not know how to fit in a bilateral gravity. She might consider every pairwise link connecting two firms. As from Figure B2, it would imply an inflation of linkages. The analyst breaks the unity of ownership linkages, while every middlemen subsidiary is technically considered as a parent company originating its own investment choice. Please note that this setup reflects the common approach of using FDI data to proxy MNE relationships. See Bénassy-Quéré et al. (2007); Stein and Daude (2007); Egger and Pfaffermayr (2004); Blonigen et al. (2003). FDI data provided by the OECD is collected following the 4th edition of the Benchmark Definition of FDI (BMD4), which primarily attributes inward FDI stocks to the immediate partner country in the transaction.
In Table B1, we report the results of our three naïve applications. The main coefficients align with the magnitude and sign we expect from these frictions, except in case a), where the coefficient on the N. of overlapping working hours is not statistically significant. In either case, findings might be irrelevant because we know that each application misses a point, and they are overall wrong because they neglect the existence of ownership chains, which are an essential ingredient in the organization of MNEs.
@@
Hines Jr and Rice (1994)’s tax
havens list
A binomial variable flagging countries identified as tax havens by Hines Jr and Rice (1994), i.e.
Andorra, Anguilla, Antigua and Barbuda, Aruba, Bahamas, Bahrain, Barbados, Belize, Bermuda,
British Virgin Islands, Cayman Islands, Cook Islands, Cyprus, Dominica, Gibraltar, Grenada,
Guernsey, Hong Kong, Ireland, Isle of Man, Jersey, Lebanon, Liberia, Liechtenstein, Luxembourg,
Malta, Marshall Islands, Mauritius, Monaco, Montserrat, Netherlands Antilles, Panama, Samoa,
San Marino, Seychelles, Singapore, St. Kitts and Nevis, St. Lucia, St. Vincent and the Grenadines,
Switzerland, Turks and Caicos Islands, Vanuatu.
Hines Jr and Rice (1994)
Hines Jr (2010)s tax havens list
A binomial variable flagging countries identified as tax havens by Hines Jr (2010), who
update the 1994 list to include Channel Islands (Guernsey and Jersey), Jordan, Macao, Maldives,
Nauru, Niue, Tonga, U.S. Virgin Islands, United Arab Emirates.
Hines Jr (2010)
@@
O 84 Public administration and defence 835 0.07% 212 0.08%
@@
U 99 Activities of extraterritorial organisations 149 0.01% 25 0.01%

### v3

@@ 3.2 Stylized facts II: hierarchical positions and geographical distances
Table 4: How distant are subsidiaries from parent companies? Complex MNEs
@@
  • Note: The table reports descriptive evidence on foreign subsidiaries controlled by multinational enterprises. The simplest MNEs with only one direct subsidiary are excluded. Columns 1, 2 and 3 on the left indicate the position in the ownership hierarchy, i.e. how many control links run from the HQs to the single subsidiary. Columns 4, 5, and 6, in the centre, count how many national borders are crossed along an ownership chain to reach a final subsidiary from the headquarters. Columns 7, 8, and 9 on the right report the geographic distance (thousands km) separating final subsidiaries from parent companies. Distance is measured between the most populated cities of hosting countries.
@@
Finally, we take a more classical approach by measuring the physical distance in kilometers between headquarters and subsidiaries, as reported in columns 7 to 9 of Table 4. These data, obtained from Conte et al. (2022) (CEPII database), represent the distance between the main cities in the countries where the headquarters and subsidiaries are located. One could hardly expect that 47.3% of the subsidiaries are located within 5,000 km of the parent company, a somewhat narrower range than expected. Beyond this threshold, the distribution has a long right tail, with subsidiaries located at increasing distances from the parent. [^15] The longest observed distance in our dataset is 19,950 km, between Paraguay and Taiwan.
@@ 4 A structural model for ownership chains
Motivated by previous empirical evidence, we present a simple model to explain location strategies along ownership chains in multinational enterprises (MNEs). Inspired by Head and Ries (2008), we assume a competitive contest in which headquarters compete to invest in new subsidiary assets worldwide while incurring monitoring costs. Specifically, in the case of ownership chains, each headquarters faces a choice: it can either directly monitor a final target subsidiary or delegate monitoring to an intermediate managerial unit - an intermediate affiliate (sometimes referred to as a middleman). These intermediate affiliates are subsidiaries that are already part of the corporate structure before the auction takes place. A subsidiary is added to an ownership chain when the headquarters delegates or self-delegates monitoring, i.e. when the cost of delegation is effectively zero. In the latter case, we have the special case of direct control.
@@
Our model abstracts from production processes between headquarters and subsidiaries. While our setup does not exclude the possibility of supply chain coordination, it also applies to conglomerates and diversified business portfolios, where monitoring remains crucial for governance and managerial control.161616 To assess whether the length of ownership chains is associated with increasing distance in the production space, we ran a regression in which the hierarchical distance of an affiliate from its parent served as the dependent variable, including, among other controls, the “production distance”, i.e. $upstreamness_{ij}$, a measure from Alfaro et al. (2019). While the coefficient was statistically significant, its magnitude was negligible. This suggests that an implausibly long ownership chain would be required for an economically meaningful difference in production stages between parent and subsidiary. In our data, these results suggest that ownership chains are primarily shaped by organizational considerations rather than sequential production optimization. Results are available upon request.
@@ 4.1 The inspection game
In this corporate governance game, profits arise from the value generated by the headquarters (Hq) and its existing perimeter of subsidiaries ($A$), [^18] The existence of a corporate perimeter before new acquisitions implies that $A=\sum_{f}a_{f}$, meaning that the total value is generated by both Hq and all existing subsidiaries, regardless of new acquisitions. as well as by the additional subsidiary ($b$), conditional on the effort exerted by the subsidiary’s manager ($e$). Headquarters can choose to monitor $q$ or trust $1-q$ the subsidiary. Similarly, the subsidiary’s manager can work ($1-p$) or shirk ($p$). If the subsidiary works, headquarters pays a wage $w$; however, if headquarters monitors and discovers shirking, the wage and the value of the subsidiary default to zero ($w=b=0$).
@@
Since corporate control operates in a global market, all Hqs play the same inspection game when evaluating subsidiaries that are available for acquisition. Consequently, competition ensures that the winning Hq is the one offering the highest valuation, determined by equation 2. It is important to note that, compared to the classical literature on delegation and monitoring in hierarchical structures (Holmstrom and Tirole, 1989), our framework relies on a number of simplifying assumptions regarding the contractual environment. First, we assume no information asymmetry between Hqs and intermediate affiliates (Mids), i.e. each party perfectly observes the other’s effort. Second, we exclude the possibility of side contracting, thus eliminating potential collusion between Mids and Subs. Finally, we treat the effort made by Mids as exogenous. In the next section, we examine how these assumptions shape Hq’s decision process and monitoring cost design.
@@ 4.2.2 The case of direct control
The simplest strategy for Hq is to retain full corporate control and not delegate monitoring to a Mid. In this case, Hq self-delegates, which means that delegation costs are effectively zero. More formally, in our model, direct control is treated as a valid location choice without delegation, represented as a $ikj$-th triplet where the intermediate jurisdiction ($k$) is null, i.e. $k=\emptyset$.
@@ 4.3.1 Locations of the chain
where $\mathcal{M}_{ikj}=1$ if Hq in country $i$ willing to invest in country $j$ chooses its monitoring unit in country $k$. F[•] is the joint cumulative distribution function, conditional on the value of $\epsilon_{ikj}$. By integrating $F[\bullet]$ with respect to the marginal probability density function $f(\epsilon_{ikj})$, we obtain the unconditional probability - the probability that country $k$ is chosen as the monitoring location for any given realization of $\epsilon_{ikj}$.
@@
where we can define $\mathcal{C}_{ij}=-ln[\sum\limits_{\ell}e^{-(\delta_{i\ell}+\delta_{\ell j})}]$ as the expected aggregate cost of locations $j$ for Hqs in country $i$, whatever the intermediate country $k$.
@@ 4.3.2 Probability of winning the auction for corporate control
The highest valuation wins the auction, i.e. the marginal probability of a Hq in country $i$ gaining control over a Sub in country $j$ is given by the probability that the highest valuation for a given Sub in $j$ comes from an Hq located in $i$, i.e. the maximum valuation in $i$ is the highest among the maxima of all competing countries. This probability is:
@@ 6 Results
Finally, the last column of Table 6 presents the coefficient of the vector of the expected cost of corporate control ($E[C_{ikj}]$), obtained from previous estimates. The coefficient obtained corresponds to $\theta$ in equation 9, and its sign and magnitude are as expected: an increase in the cost index reduces the number of subsidiaries in country $j$ controlled by parents in country $i$. Figure 8 shows the distribution of the expected inspection cost for each origin $i$ and destination $j$, which is a useful byproduct of our structural model.
Figure 8: Distribution of the inspection cost
  • Note: The figure shows the distribution of the expected cost of corporate control estimated after Eq. 8. The value can be negative or positive depending on the prevalence of gravity components (overlapping working hours, geographical distance, regional trade agreements, common language, colony dependence, legal origins) that proxy frictions with a negative or positive coefficient.
@@ 7 Robustness and sensitivity checks
  • Note: Ratios $CT_{k}/CT_{i}$ and $CT_{j}/CT_{k}$ control for relative advantage in corporate taxation. Ratios $LC_{k}/LC_{i}$ and $LC_{j}/LC_{k}$ control for the relative advantage in labor costs. Given the presence of zeros in the case of direct control, we scale the distances in km by using the hyperbolic sine transformation ($ln(x+\sqrt{x^{2}+1})$), which allows approximating to the natural logarithm while retaining the zeros (Bellemare and Wichman, 2020). Standard errors are double-clustered by origin and destination in parentheses (*** p$<$0.01, ** p$<$0.05, * p$<$0.1).
@@ 8 Conclusions
Building on these empirical patterns, we propose a location choice model in which parent firms compete in a global auction to acquire new subsidiaries. In formulating their bids, parents weigh the option of delegating monitoring responsibilities to an intermediate affiliate in a third country. The model yields two testable equations: the first is a trilateral equation that determines whether monitoring is delegated to an intermediate affiliate and where it is located, while the second is a bilateral equation that predicts the location of newly acquired final subsidiaries.
@@
In this Appendix, we provide an understanding of how raw shareholding data are used to extract corporate trees developed by MNEs. For further details, we refer to Rungi et al. (2017). An MNE is made of a unique parent company and its subsidiaries. If we look at their ownership data, they represent a corporate network in the form of a hierarchical tree, where nodes are legally autonomous companies. Please consider the simplified fictional corporate tree in Figure A1. On top of the hierarchy, we find the parent company, which exerts control over subsidiaries by means of direct and indirect shareholding links. In principle, an MNE can develop much more complex patterns than the ones depicted in Figure A1, as it can include cross-holdings when shareholding relationships run two-way between companies. They can also entail ownership cycles, i.e. when the firm holds a share of its equity stakes. Given a fragmentation of shareholding, companies can indirectly connect through multiple sequences of ownership links. The solution is to start with a full ownership matrix of all firms and their shareholders in the data. Thus, a control relationship is detected iteratively on a transformation of the ownership matrix. In line with international accounting standards (OECD, 2005; UNCTAD, 2009; Eurostat, 2007), one can assume that corporate control is present when there is an absolute majority of equity stakes. Yet, one cannot exclude that lower thresholds still allow the parent company the ability to influence and contribute to the decision-making of another company.
@@
In Figure A1, we easily attribute direct control of Subsidiary 1 to the Parent because the latter holds an absolute majority of voting rights in the first. Besides this most elementary case of direct control, there are two other ways for the Parent to exert control: (i) by transitivity, when successive majority links form an ownership chain, and subsidiaries align on it, as in the case of Subsidiary 3; (ii) by consolidation, when equity is fragmented, no direct majority path can be detected from the Parent, and yet other subsidiaries hold minority links in a firm, like Subsidiary 4, which can be consolidated and allow the unique Parent to finally reach out to the management of Subsidiary 4.
Ownership chains. Figure A1 shows an MNE that has two ownership chains. One starts from the Parent and reaches out to Subsidiary 3 through Subsidiary 1. A second one starts from the Parent and reaches out to Subsidiary 4 thanks to either Subsidiary 1 or Subsidiary 3. To solve uncertainty, we will assume that the final subsidiary is on the shortest ownership chain. If ownership chains are of equal length, then we discriminate on shares. When shares are equal, we discriminate by firm size, giving priority to the middleman that is bigger.
@@
In the following paragraphs, we introduce three naïve unstructured applications of gravity models to observe correlations, which are informative for the setup of our theoretical model, and to discuss what the analyst misses when not considering information on ownership chains. We argue that this section is important to identify empirical problems from previous literature. In each case, we test the following general Pseudo-Poisson specification at the country level:
@@
where $N_{ij}$ counts the number of companies in country $j$ controlled by companies in country $i$. We include bilateral frictions in matrix $\mathbf{X}_{ij}$. Notably, besides classical gravity controls sourced from the CEPII Gravity dataset (Conte et al., 2022) $wh_{ij}$, we include a measure of overlapping working hours between countries $i$ and $j$ as a proxy of the common working schedule across time zones. See further details in Section 3. Finally, $\gamma_{i}$ and $\gamma_{j}$ capture origin and destination country fixed effects. [^22] Please note that our information on MNEs’ perimeters is cross-sectional. Therefore, we can focus on the between-variation and cannot observe the within-variation.
At this point, we are ready with the empirical model, but how do we treat ownership chains? We will show that each of the three following cases is wrong. Or, we could say that they are reduced forms of the real model. Let us introduce the issue starting from a fictional sample, including two kinds of MNEs most common in our data set, whose fictional structures are represented in Figure B1.
@@
  • Note: This fictional dataset includes a simple and a complex MNE. Each node represents a company, and each link represents a company’s controlling voting rights in another company.
@@
  • Notes: The figure shows three fictional cases of ownership chains included in bilateral gravity models. The legends on the right report the (directed) links one would include in each application.
More specifically, in Figure B1, we draw one simple MNE controlling only one subsidiary, which we know is the most common case, although we also know that, on aggregate, these MNES have a small economic weight (see Figures 3 and 4 ). A second complex MNE on the right controls seven subsidiaries organized along four ownership chains: i) A $\rightarrow$ B; ii) A $\rightarrow$ C $\rightarrow$ E $\rightarrow$ H; iii) A $\rightarrow$ D $\rightarrow$ F; iv) A $\rightarrow$ D $\rightarrow$ G.
@@
Let us first assume that linkages are not disclosed, but the set of subsidiaries is known. In this case, the analyst will count the parents in country $i$ and the subsidiaries in country $j$. This is the case (a) in the fictional sample redrawn in Figure B2. We guess this is the case that most scholars have in mind when testing gravity models with firm-level data on MNEs. See, for example, Mutti and Grubert (2004), which reports an empirical analysis using data on the location choices of US parent companies.
A second alternative is that ownership chains are only partially disclosed because the statistician has problems collecting them. S/he falsely attributes the ultimate control of a subsidiary to its immediately controlling corporate shareholder. Hence, our fictional complex MNE will appear fragmented into three smaller pieces. The information included in Eq. 10 is false. This is the second case in Figure B2, where the list of links on the right does not include $C\rightarrow E$. Indeed, full disclosure of the corporate structure of MNEs is a long-standing problem for national and international statistics offices. The necessity to identify ultimate controlling entities is required by several guidelines (OECD, 2005; UNCTAD, 2009; IFRS, 2011), seldom complied with.
A third alternative is that the analyst knows the entire ownership chain but does not know how to fit in a bilateral gravity. S/he might consider every pairwise link connecting two firms. As from Figure B2, it would imply an inflation of linkages. The analyst breaks the unity of ownership linkages, while every middlemen subsidiary is technically considered as a parent company originating its own investment choice. Please note that this setup reflects the common approach of using FDI data to proxy MNE relationships. See Bénassy-Quéré et al. (2007); Stein and Daude (2007); Egger and Pfaffermayr (2004); Blonigen et al. (2003). FDI data provided by the OECD is collected following the 4th edition of the Benchmark Definition of FDI (BMD4), which primarily attributes inward FDI stocks to the immediate partner country in the transaction.
In Table B1, we report the results of our three naïve applications. The main coefficients are consistent with the magnitude and sign we would expect from these frictions, except in case a), where the coefficient on the N. of overlapping working hours is not statistically significant. In either case, findings might be irrelevant because we know that each application misses a point, and they are wrong overall because they neglect the existence of ownership chains, which are an essential ingredient in the organization of MNEs.
@@
O 84 Public administration and defense 835 0.07% 212 0.08%
@@
U 99 Activities of extraterritorial organizations 149 0.01% 25 0.01%